# May the Feedback Be with You! Unlocking the Power of Feedback-Driven Deep Learning Framework Fuzzing via LLMs

### v3

@@ 2.2. LLM-based Fuzzing
Figure 5. Default prompt template of FUEL
3. FUEL Design
Figure 4 depicts the end-to-end workflow of FUEL, integrating a $\mathcal{M_{A}}$ with a feedback-aware simulated annealing algorithm to utilize feedback information in a fully automated manner and subsequently combining analysis results to guide $\mathcal{M_{G}}$ to generate diverse and valid DL models. We show the details in the rest of this section.
3.1. Prompting LLM for Test Generation
To accommodate these two scenarios, we design specialized default generation prompts that ensure continuous test generation even in scenarios where feedback is unavailable or repair stagnation occurs. The transition between feedback-driven and default generation is algorithmically controlled to optimize exploration-exploitation balance throughout the whole fuzzing campaign. Figure 5(a) shows a few-shot example in the default generation prompt. Firstly, it lists some operators selected by the heuristic search algorithm (discussed in § 3.4). Then this example goes with the instruction “Please generate a valid PyTorch model with selected operators”. Next, it provides a DL model consisting of two components: model definition and input tensors (introduced in § 2.1). Finally, with such few-shot examples, $\mathcal{M_{G}}$ generates tests with the prompt template shown in Figure 5(b). [LIBRARY] denotes the specific framework (PyTorch or TensorFlow), [SELECTED_OPS] is filled by results of heuristic search, and [PLACEHOLDER...] is the DL model to be generated by $\mathcal{M_{G}}$.
3.2. Differential Testing and Oracles
Differential Testing. Differential testing has been proven to be an effective way to detect bugs in DL frameworks (Liu et al., 2023a). The DL model generated by LLM is executed on different backends to identify bugs. In this work, we focus on the compilation feature of DL frameworks and design the cross-check mode on two different backends as follows.
@@ 3.2. Differential Testing and Oracles
  • Compiler backend has been the most important feature of DL frameworks since the version 2.x era, pushing the performance of DL frameworks to new heights. As an emerging technology, the compiler backend is still in rapid development, and has become an important backend to compile DL models (especially LLMs) (Yang et al., 2024). PyTorch Inductor and TensorFlow XLA are compiler backends for PyTorch and TensorFlow, respectively.
@@
  • Behavior inconsistency (program crashes). In addition to numerical inconsistency, aligning the behavior between the eager backend and the compiler backend is also very crucial. Both eager backend and compiler backend should behave the same on the same test inputs (i.e., both throw errors or pass tests). Otherwise, we consider this a potential bug (i.e., one throws errors and the other one does not).
Execution and cross-checking on the two backends can be classified into three distinct states (Figure 1): (i) Cross-checking results violate the oracles, indicating a potential bug in DL frameworks, then we collect bug reports. (ii) If a DL model crashes on both backends, it is an invalid model and we collect exception logs. (iii) Only when execution results of both backends do not violate oracles or crash, it is a bug-free and valid model, and code coverage is collected.
3.3. Prompting LLM for Analysis Summary
Figure 6. Analysis prompt templates of FUEL
@@ 3.3. Prompting LLM for Analysis Summary
3.4. Heuristic Search for Operators Selection
@@
While LLM-based approaches can benefit from sufficient analysis summaries, they tend to use some common operators (e.g., Conv2d, ReLU, and Linear) for test generation (Deng et al., 2024b), which significantly limits model diversity (discussed in Limitation #4). This limitation introduces the implementation of RAG (Gao et al., 2023) for $\mathcal{M_{G}}$, a logical and natural enhancement that prioritizes less prevalent operators in test generation.
@@
Unfortunately, operator selection is a challenging NP-hard problem, stemming from the combinatorial explosion inherent in DL model construction: While DL frameworks typically offer a finite operator set (numbering in the thousands), the unbounded combinations of these operators create an intractably large retrieval space. Notably, we identify that feedback signals can serve as indicators of intrinsic value, enabling the iterative optimization of operator selection through value-guided search. The idea behind employing heuristic search lies in reformulating the RAG task as a combinatorial optimization problem within a constrained space: At each iteration, the algorithm selects an optimal subset of operators from the finite collection that maximizes a predefined value function. This value function, dynamically determined by accumulated feedback signals, quantifies the strategic importance of operators for subsequent iterative generation.
@@
Inspired by the above idea, we adopt simulated annealing (Van Laarhoven et al., 1987) for operator selection in FUEL, a balance between overhead and performance. High-performance search algorithms, such as Monte Carlo Tree Search, incur prohibitive temporal overhead during incremental tree construction in iterative fuzzing cycles. Conversely, low-overhead search algorithms (e.g., random sampling) fail to identify high-value operators, compromising search performance. Simulated annealing emerges as an offline planning algorithm that efficiently approximates global optima (regulated through cooling rate $\gamma$) within constrained time budgets. For operator valuation, we design three metrics of operators: (i) Number of uses; (ii) Exception-triggering count; (iii) Number of new triggered line coverage. The valuation function follows an inverse relationship with the first two factors (i.e., frequency and exceptions), reflecting our emphasis on exploring underused operators. Conversely, it maintains direct proportionality to new code coverage, prioritizing operators that can trigger new code coverage. We define the final valuation function as follows: $V_{\alpha,\beta}(x,y,z)=\frac{\alpha}{\alpha+x}+\frac{\alpha}{e^{y}}+\beta-e^{-\frac{z}{100}}$, where $x$, $y$, and $z$ denote for number of use, exception-triggering count, and number of new triggered line coverage, respectively. The hyper-parameters $\alpha$ and $\beta$ are both initialized at $0.5$. This configuration ensures algorithmic stability before the fuzzing campaign, where $x$, $y$, and $z$ remain zero-initialized. Under this initialization setting, the value for all operators reduces to: $V_{\alpha,\beta}(0,0,0)=\alpha+\beta$. Substituting the predefined hyperparameters yields a uniform initial value $V_{0.5,0.5}(0,0,0)=1.0$, ensuring fair operator prioritization at the beginning of fuzzing.
1
2
3
4
@@
6  if exceptionFlag ==1==1 then
7     selectedOps $\leftarrow$ lastOps
8     return selectedOps
9 if isNotFirstIteration then
10     for $\textnormal{{Op}}\in\textnormal{{lastOps}}$ do
11        $\textnormal{{Op}}_{used\_times}\leftarrow\textnormal{{Op}}_{used\_times}+1$
12        $\textnormal{{Op}}_{exp\_count}\leftarrow\textnormal{{Op}}_{exp\_count}+$ exceptionFlag
@@
14
15
16 selectedOps $\leftarrow$ SimulatedAnnealing(opSet)
17  return selectedOps
18
19Function SimulatedAnnealing(opSet):
20  $T,N_{s},\gamma$ $\leftarrow$ $100,10,0.99$
21  $k$ $\leftarrow$ RandomNumber($1,3$)
22  $S_{cur}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
23  while T $\geq$ $T_{min}$ do
24     $n_{s}$ $\leftarrow$ 0
25     while $n_{s}$ ¡ $N_{s}$ do
26        $S_{new}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
@@
29           $S_{cur}$ $\leftarrow$ $S_{new}$
30       $n_{s}\leftarrow n_{s}+1$
31
32    $T\leftarrow T\cdot\gamma$
33 return $S_{cur}$
@@ 32    $T\leftarrow T\cdot\gamma$
Algorithm 1 presents the details of our heuristic search: Feedback-Aware Simulated Annealing (FASA). The algorithm begins by verifying whether an exception was triggered in the previous iteration (Line 1). If true, FASA returns the previously selected operator sequence to provide a repair opportunity for FUEL (Lines 1-1). Subsequently, operator values are updated through the integration of three distinct feedback signals (Lines 1-1). The optimization process employs a simulated annealing framework that returns the selected operators (Lines 1-1), initialized with standard temperature parameters and cooling schedules (Lines 1-1). During each annealing cycle (Line 1), $N_{s}$ iterations are performed to evaluate candidate operator sequences $S_{new}=[op_{1},op_{2},...,op_{k}]$ (Lines 1-1), where the composite fitness value function $F$ is calculated as the average of individual operator values ($F=\frac{1}{k}\sum_{i=1}^{k}V_{\alpha,\beta}(x_{i},y_{i},z_{i})$). While sequences with $\Delta F\geq 0$ are typically accepted, the Metropolis criterion probabilistically permits inferior solutions (with probability $e^{\Delta F/T}$) to escape local optima (Lines 1-1). The temperature parameter undergoes reduction according to a predefined cooling schedule until reaching termination thresholds (Lines 1-1). Finally, FASA returns a selected operator sequence in the current fuzzing iteration (Line 1).
3.5. Feedback-guided Prompting for Test Generation
Figure 7. Feedback-guided prompt template of FUEL
@@ 4. Implementation
Operator collection from API document. As discussed in § 3, an operator set is needed for FASA. In this work, we collect core operators from two DL frameworks. Specifically, we implement an HTML crawler to collect operators from PyTorch with API prefixes torch, torch.nn, and torch.nn.functional as they are core operators (PyTorch, 2021) in PyTorch 2.x. For TensorFlow, we collect operators that XLA supports from its document (TensorFlow, 2025) as we focus on the compilation feature of the DL framework in this work. Although we currently only consider the most important subset of all operators, we can easily modify the operator set to suit specific test scenarios (e.g., considering gradient-related operators to test the gradient feature of DL frameworks).
@@
Generation and analysis LLMs. While closed-source models (e.g., GPT-4o (GPT-4o, 2023)) demonstrate amazing performance, their costs are substantial through API service. In contrast, open-source models enable local deployment but typically exhibit comparatively inferior performance. In our approach, we use DeepSeek-V3 (DeepSeek-AI, 2024) as the $\mathcal{M_{A}}$ for feedback analysis, leveraging its demonstrated superiority in most benchmarks related to text summary. Meanwhile, we select Qwen-2.5-Coder-32B (Hui et al., 2024) as $\mathcal{M_{G}}$ for test generation due to its leading performance on code generation benchmarks (EvalPlus, 2025). The temperature of the Qwen model is set to one that aligns with the mechanism of fuzzing (i.e., generating random and different test cases each time). This model selection strategy carefully balances the trade-off between performance and costs. Specifically: (i) For local deployment, we use vLLM (vLLM-project, 2025), an efficient LLM inference engine with optimized memory management, to host open-source local models; (ii) For API services, we adopt a Python package openai (OpenAI, 2025) that maintains compatibility with mainstream model API services. Notably, our artifact supports flexible integration of diverse LLMs through different invocation ways (i.e., local deployment and API services), allowing dynamic selection between generation and analysis models based on specific requirements and budgets.
5. Evaluation
5.1. Experimental Setup
@@
Systems under test. In this work, we focus on two popular DL frameworks (PyTorch (PyTorch, 2018) and TensorFlow (TensorFlow, 2020)). The statistics of two DL frameworks are shown in Table 1.
Table 1. Statistics of the SUTs
Lines of Code Test Language Tested Version
PyTorch 322,839 Python Nightly20250214
TensorFlow 285,406 Python Nightly20250228
@@
Baselines. In this work, we select two SOTA fuzzers for DL framework testing (i.e., TitanFuzz (Deng et al., 2023) and WhiteFox (Yang et al., 2024)). Notably, traditional fuzzers are not included because WhiteFox has demonstrated that LLM-based fuzzers can outperform the SOTA traditional fuzzer (e.g., NNSmith (Liu et al., 2023a)) on both code coverage and bug detection ability.
@@
Metrics. We consider two types of metrics. (i) Code coverage: We use a Python library (namely Coverage.py (cov, 2022)) to collect line code coverage. (ii) Number of detected bugs: Notice that we also record the number of detected bugs. We report our detected bugs to the corresponding repository issues on GitHub. We record the issue as “Confirmed” only when the developer can reproduce it and as “Fixed” when a pull request (PR) linked to the issue is merged into the main branch.
5.2. Comparison with SOTA baselines
5.2.1. Default fuzzing configuration
Figure 8 demonstrates the performance of FUEL against two SOTA baselines (TitanFuzz and WhiteFox), across both on PyTorch and TensorFlow. Over 1,000 iterations, FUEL achieves superior code coverage compared to both baselines, outperforming the second-best baseline TitanFuzz by 9.15% and 14.70% on PyTorch and TensorFlow, respectively. Notably, this enhanced code coverage is obtained despite FUEL employing fewer operators, indicating its effectiveness in combining different operators. This performance advantage may benefit from the model-level fuzzing paradigm of FUEL that comprehensively explores diverse model structures compared to API-level fuzzers. In contrast, WhiteFox exhibits relatively lower coverage, potentially attributed to its optimization-triggering pattern specialization. While it is effective for testing specific compiler optimizations, this feature inadvertently restricts model diversity through excessive template matching. Experimental results support our hypothesis that utilizing coverage information to guide test generation proves more effective than the API-level fuzzing technique or optimization-targeted approach.

### v4

@@ 3.1 Prompting LLM for Test Generation
Figure 5: Default prompt template of FUEL
3.2 Differential Testing and Oracles
Differential Testing. Differential testing has been proven to be an effective way to detect bugs in DL frameworks [liu2023nnsmith]. The DL model generated by LLM is executed on different backends to identify bugs. In this work, we focus on the compilation feature of DL frameworks and design a cross-check mode on two different backends as follows.
@@ 3.2 Differential Testing and Oracles
  • Compiler backend has been the most important feature of DL frameworks since the version 2.x era, pushing the performance of DL frameworks to new heights. As an emerging technology, the compiler backend is still in rapid development, and has become an important backend to compile DL models (especially LLMs) [whitefox]. PyTorch Inductor and TensorFlow XLA are compiler backends for PyTorch and TensorFlow, respectively.
@@
  • Behavior inconsistency (program crashes). In addition to numerical inconsistency, aligning the behavior between the eager backend and the compiler backend is also very crucial. Both eager backend and compiler backend should behave the same on the same test inputs (i.e., both throw errors or pass tests). Otherwise, we consider this a potential bug (i.e., one throws errors, and the other one does not).
Execution and cross-checking on the two backends can be classified into three distinct states (Figure 1): (i) Cross-checking results violate the oracles, indicating a potential bug in DL frameworks, then we collect bug reports. (ii) If a model crashes on both backends, it is an invalid model, and we collect exception logs. (iii) Only when execution results of both backends do not violate oracles or crash, it is a bug-free and valid model, and code coverage is collected.
3.3 Prompting LLM for Analysis Summary
Figure 6: Analysis prompt templates of FUEL
@@ 3.3 Prompting LLM for Analysis Summary
3.4 Heuristic Search for Operator Selection
@@
While LLM-based approaches can benefit from sufficient analysis summaries, they tend to use some common and well-tested operators (e.g., Conv2d, ReLU, and Linear) for test generation [fuzzgpt], which significantly limits model diversity (discussed in Limitation #4). This limitation introduces the implementation of providing additional context information for $\mathcal{M_{G}}$, a logical and natural enhancement that prioritizes less prevalent operators in test generation.
@@
Unfortunately, operator selection is a challenging NP-hard problem, stemming from the combinatorial explosion inherent in DL model construction: While DL frameworks typically offer a finite operator set (numbering in the thousands), the unbounded combinations of these operators create an intractably large retrieval space. Notably, we identify that feedback signals can serve as indicators of intrinsic value, enabling the iterative optimization of operator selection through value-guided search. The idea behind employing heuristic search lies in reformulating the operator selection task as a combinatorial optimization problem within a constrained space: At each iteration, the algorithm selects an optimal subset of operators from the pre-collected API documents that maximizes a predefined value function. This value function, dynamically determined by accumulated feedback signals, quantifies the strategic importance of operators for subsequent iterative generation.
@@
Inspired by the above idea, we adopt simulated annealing [van1987simulated] for operator selection in FUEL, a balance between overhead and performance. High-performance search algorithms, such as Monte Carlo Tree Search, incur prohibitive temporal overhead during incremental tree construction in iterative fuzzing cycles. Conversely, low-overhead search algorithms (e.g., random sampling) fail to identify high-value operators, compromising search performance. Simulated annealing emerges as an offline planning algorithm that efficiently approximates global optima (regulated through cooling rate $\gamma$) within constrained time budgets. For operator valuation, we design three metrics of operators: (i) Number of uses; (ii) Exception-triggering count; (iii) Number of new triggered line coverage. The valuation function follows an inverse relationship with the first two factors (i.e., frequency and exceptions), reflecting our emphasis on exploring underused operators. Conversely, it maintains direct proportionality to new code coverage, prioritizing operators that can trigger new code coverage. This strategy effectively prevents the fuzzer from being trapped in coverage plateaus caused by over-testing common operators. We define the final valuation function as follows: $V_{\alpha,\beta}(x,y,z)=\frac{\alpha}{\alpha+x}+\frac{\alpha}{e^{y}}+\beta-e^{-\frac{z}{100}}$, where $x$, $y$, and $z$ denote for number of use, exception-triggering count, and number of new triggered line coverage, respectively. The hyper-parameters $\alpha$ and $\beta$ are both initialized at $0.5$. This configuration ensures algorithmic stability before the fuzzing campaign, where $x$, $y$, and $z$ remain zero-initialized. Under this initialization setting, the value for all operators reduces to: $V_{\alpha,\beta}(0,0,0)=\alpha+\beta$. Substituting the predefined hyperparameters yields a uniform initial value $V_{0.5,0.5}(0,0,0)=1.0$, ensuring fair operator prioritization at the beginning of fuzzing.
Algorithm 1 presents the details of our Feedback-Aware Simulated Annealing (FASA). The algorithm begins by verifying whether an exception was triggered in the previous iteration (Line 1). If true, FASA returns the previously selected operator sequence to provide a repair opportunity for FUEL (Lines 1-1). Subsequently, operator values are updated through the integration of three distinct feedback signals (Lines 1-1). The optimization process employs a simulated annealing framework that returns the selected operators (Lines 1-1), initialized with standard temperature parameters and cooling schedules (Lines 1-1). During each annealing cycle (Line 1), $N_{s}$ iterations are performed to evaluate candidate operator sequences $S_{new}=[op_{1},op_{2},...,op_{k}]$ (Lines 1-1), where the composite fitness value function $F$ is calculated as the average of individual operator values ($F=\frac{1}{k}\sum_{i=1}^{k}V_{\alpha,\beta}(x_{i},y_{i},z_{i})$). While sequences with $\Delta F\geq 0$ are typically accepted, the Metropolis criterion probabilistically permits inferior solutions (with probability $e^{\Delta F/T}$) to escape local optima (Lines 1-1). The temperature parameter undergoes reduction according to a predefined cooling schedule until reaching termination thresholds (Lines 1-1). Finally, FASA returns a selected operator sequence in the current fuzzing iteration (Line 1).
@@
 if exceptionFlag ==1==1 then
    selectedOps $\leftarrow$ lastOps
    return selectedOps
   end if
 if isNotFirstIteration then
    for $\textnormal{{Op}}\in\textnormal{{lastOps}}$ do
       $\textnormal{{Op}}_{used\_times}\leftarrow\textnormal{{Op}}_{used\_times}+1$
       $\textnormal{{Op}}_{exp\_count}\leftarrow\textnormal{{Op}}_{exp\_count}+$ exceptionFlag
@@
      end for
   end if
 selectedOps $\leftarrow$ SimulatedAnnealing(opSet)
 return selectedOps
Function SimulatedAnnealing(opSet):
 $T,N_{s},\gamma$ $\leftarrow$ $100,10,0.99$
 $k$ $\leftarrow$ RandomNumber($1,3$)
 $S_{cur}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
 while T $\geq$ $T_{min}$ do
    $n_{s}$ $\leftarrow$ 0
    while $n_{s}$ < $N_{s}$ do
       $S_{new}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
@@
          $S_{cur}$ $\leftarrow$ $S_{new}$
         end if
       $n_{s}\leftarrow n_{s}+1$
      end while
    $T\leftarrow T\cdot\gamma$
   end while
 return $S_{cur}$
@@
3.5 Feedback-guided Prompting for Test Generation
Figure 7: Feedback-guided prompt template of FUEL
@@ 4 Implementation
Operator collection from API document. As discussed in § 2.1, an operator set is needed for FASA. In this work, we collect core operators from two DL frameworks. Specifically, we implement an HTML crawler to collect operators from PyTorch with API prefixes torch, torch.nn, and torch.nn.functional as they are core operators [ptop] in PyTorch 2.x. For TensorFlow, we collect operators that XLA supports from its document [tfxlaops] as we focus on the compilation feature of the DL framework in this work. Although we currently only consider the most important subset of all operators, we can easily modify the operator set to suit specific test scenarios (e.g., considering gradient-related operators to test the gradient feature of DL frameworks).
@@
Generation and analysis LLMs. While closed-source models (e.g., GPT-4o [gpt-4o]) demonstrate amazing performance, their costs are substantial through the API service. In contrast, open-source models enable local deployment but typically exhibit comparatively inferior performance. In our approach, we use DeepSeek-V3 [deepseekai2024deepseekv3] as the $\mathcal{M_{A}}$ for feedback analysis, leveraging its demonstrated superiority in most benchmarks related to text summary. Meanwhile, we select Qwen-2.5-Coder-32B [hui2024qwen2.5-coder] as $\mathcal{M_{G}}$ for test generation due to its leading performance on code generation benchmarks [evalplus]. The temperature of the Qwen model is set to one that aligns with the mechanism of fuzzing (i.e., generating random and different test cases each time). This model selection strategy carefully balances the trade-off between performance and costs. Specifically: (i) For local deployment, we use vLLM [vllm], an efficient LLM inference engine with optimized memory management, to host open-source local models; (ii) For API services, we adopt a Python package openai [openai-python] that maintains compatibility with mainstream model API services. Notably, our artifact supports flexible integration of diverse LLMs through different invocation ways (i.e., local deployment and API services), allowing dynamic selection between generation and analysis models based on specific requirements and budgets.
In-Context Learning Prompting. Prompt engineering plays a pivotal role in LLM performance [sahoo2024promptsurvey]. Since FUEL integrates two distinct LLM agents ($\mathcal{M_{A}}$ and $\mathcal{M_{G}}$), we design specialized prompts for each, employing lightweight few-shot In-Context Learning (ICL) [brown2020gpt3]. To balance manual effort with system extensibility, we curate a concise set of few-shot examples that comprehensively cover the spectrum of feedback categories defined in our design: (i) Coverage Information, distinguishing whether triggering new coverage; (ii) Bug Reports, encompassing both single-backend crashes and numerical inconsistencies; (iii) Exception Logs, representing invalid executions where crashes occur across both backends. To construct high-quality examples, we collected representative DL programs from framework documentation and historical GitHub issues. These programs were executed to capture real-world <program, feedback> pairs. We then manually crafted the corresponding [ANALYSIS SUMMARY] to form ground-truth triplets. These triplets serve as few-shot examples to teach $\mathcal{M_{A}}$ how to distill feedback, and subsequently guide $\mathcal{M_{G}}$ in interpreting these summaries for test generation. Although we opt for this lightweight prompting strategy over complex paradigms (e.g., Chain-of-Thought [wei2022cot]) for efficiency, it demonstrates sufficient effectiveness in our evaluation. Notably, such implementation significantly reduces the manual burden compared to traditional fuzzing techniques [liu2023nnsmith, wang2020lemon].
5 Evaluation
5.1 Research Questions
We design the following research questions (RQs) in our evaluation:
  • RQ1: Can FUEL outperform the state-of-the-art (SOTA) baselines on coverage and bug detection?
  • RQ2: Can FUEL detect real-world new bugs?
  • RQ3: How does each component contribute to FUEL?
5.2 Experimental Setup
@@
Systems under test. In this work, we focus on two most popular DL frameworks (PyTorch [PyTorch] and TensorFlow [Tensorflow]). The statistics of two DL frameworks are shown in Table 1. In the following tables, we use PT and TF to refer to PyTorch and TensorFlow, respectively.
Table 1: Statistics of the SUTs
SUT Lines of Code Test Language Tested Version
PT 322,839 Python Nightly20250214
TF 285,406 Python Nightly20250228
@@
Baselines. We select SOTA DL framework testing baselines from two categories (LLM-based fuzzers [titanfuzz, whitefox] and traditional fuzzers [liu2023nnsmith, liu2023neuri]). To ensure a fair comparison, we use the same oracle and differential testing (as detailed in § 3.2) for baselines. We introduce the characteristics of these baselines as follows.
  • TitanFuzz [titanfuzz] is the first technique to directly leverage LLMs for fuzzing DL libraries, capitalizing on their implicit knowledge of language syntax and API constraints. It employs a generative LLM to produce seed programs and an infilling LLM to populate valid API parameters. Finally, it executes generated APIs and compares outputs across different hardware to detect bugs.
  • WhiteFox [whitefox] is an LLM-based white-box fuzzer that leverages compiler optimization source code to guide test generation. It uses an analysis LLM to extract optimization patterns as requirements, which are then used by a generation LLM to synthesize targeted test cases.
  • NNSmith [liu2023nnsmith] targets DL compilers by utilizing an SMT solver with manually defined operator specifications to generate diverse and valid computation graphs. Then it further employs a gradient-based search to find inputs that prevent floating-point exceptions, ensuring valid comparisons during differential testing.
  • NeuRI [liu2023neuri] tackles the challenge of defining complex operator constraints by employing inductive program synthesis. It infers these constraints automatically from collected valid and invalid API traces and subsequently utilizes a hybrid model generation strategy (combining symbolic and concrete execution) to synthesize diverse and valid DL models capable of supporting hundreds of operator types.
@@
Metrics. We consider two types of metrics. (i) Code coverage: We use a Python library (namely Coverage.py [coverage-py]) to collect line code coverage. (ii) Number of detected bugs: Notice that we also record the number of detected bugs. We report our detected bugs to the corresponding repository issues on GitHub. We record the issue as “Confirmed” only when the developer can reproduce it and as “Fixed” when a pull request (PR) linked to the issue is merged into the main branch.
5.3 RQ1: Comparison with SOTA baselines
5.3.1 Default fuzzing configuration
Figure 8 demonstrates the performance of FUEL against four SOTA baselines, on both PyTorch and TensorFlow. Over 1,000 iterations, FUEL achieves superior code coverage compared to all baselines, outperforming the second-best baseline NeuRI by 4.48% and 9.14% on PyTorch and TensorFlow, respectively. Among the baselines, NeuRI demonstrates the most competitive performance. Although NeuRI is also a model-level fuzzer capable of synthesizing valid models via inductive rule inference, it is fundamentally limited by its reliance on random operator combinations. This stochastic nature prevents it from systematically exploring the rich and deep interaction space between operators. In contrast, FUEL utilizes feedback-driven guidance and heuristic search to strategically combine operators, thereby unlocking harder-to-reach code paths that random combinations miss. Other baselines show distinct limitations. For instance, WhiteFox exhibits relatively lower coverage, potentially attributed to its optimization-triggering pattern specialization. While effective for testing specific compiler optimizations, this feature inadvertently restricts model diversity through excessive template matching. These experimental results support our hypothesis that utilizing coverage information to guide test generation is more effective than LLM-based fuzzers and traditional fuzzers, achieving greater coverage across 1,000 iterations.